# Atypical generic directions in Teichmüller space

### v1

@@ 4. Combinatorics of the slit curves
The proof involves an iterated argument which a blend of some results connecting short curves and subsurface projections due to Rafi [Raf05, Raf14] and Modami–Rafi [MR25], and so-called “passing-up” techniques for producing large subsurface projections satisfying certain properties due to Durham [Dur23].
@@
We note one of the main technical difficulties here is that the slit curves $\zeta_{k}$ need not correspond to boundary curves of subsurfaces with large projections for $\mu_{-},\mu_{+}$. In particular, knowing that they are pairwise filling (Proposition 2) alone is not enough to show that they spread out in $\mathcal{C}(S)$ at a uniform rate.
@@
By Proposition 2, we may assume that the slit curves $\sigma_{i}$ pairwise fill $S$. We fix a subsurface projection threshold constant $L_{0}=L_{0}(S)>0$ to be sufficiently large to make the arguments below work.
For sufficiently large $L_{0}=L_{0}(S)>0$, we can fix $\epsilon=\epsilon(L_{0},S)>0$ to be small enough so that if $\zeta$ is a simple closed curve on $S$ so that $l_{\gamma(t)}(\zeta)<\epsilon$ for some time $t$, then work of Modami–Rafi111[MR25] provides a quantitative version of [Raf05, Theorem 6.1], namely that increasingly short curves produce increasingly large subsurface projections. [MR25] provides a subsurface $V\subset S-\zeta$ (with possibly $V$ the annulus with core $\zeta$) so that $diam_{\mathcal{C}(V)}(\gamma)>L_{0}$.
@@
Consider the sequence of slit curves $\zeta_{k}$. By item (3) of Proposition 1, there exists $K_{0}=K_{0}(\epsilon)>0$ so that if $k>K_{0}$, then $l_{\gamma(t_{k})}(\zeta_{k})<\epsilon$. For each $k>K_{0}$, let $V_{k}$ denote the $L_{0}$-large subsurface in the complement $S-\zeta_{k}$ provided by [Raf05, Theorem 6.1]. We denote the collection of these subsurfaces $V_{k}$ by $\mathcal{V}$.
Our goal is to show that the $\zeta_{k}$ spread out in $\mathcal{C}(S)$ uniformly quickly. Using the $V_{k}$ as surrogates, the following claim forces the boundary curves of any sufficiently large subcollection to spread out in some subsurface, with extra control over where the slit curves lie:
@@
For any $L_{1}>L_{0}$, there exists $N_{1}=N_{1}(S,L_{0},L_{1})>0$ so that for any collection $\mathcal{U}\subset\mathcal{V}$ with #$\#\mathcal{U}\geq N_{1}$, there exists a subsurface $W$ and $V_{j},V_{j},V_{k},V_{l}\subset W$ with $V_{i},V_{j},V_{k},V_{l}\in\mathcal{U}$ so that
@@
    All pairwise distances in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+},\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ are larger than $\frac{L_{1}}{100}$, and $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ appear in that order along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$ (see Figure 2).
@@
[…]
Figure 2. Using [Dur23, Proposition 4.7], we can arrange for the boundary curves $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$, and hence $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$, appear in order along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$.
@@
Items (1) and (2) are essentially an immediate consequence of [Dur23, Proposition 4.7], which not only provides a large domain $W$ containing some of the $V_{k},\dots,V_{k+N_{1}}$ used to produce it, but forces some of their boundary curves to roughly evenly distribute along the geodesic in $\mathcal{C}(W)$ between $\pi_{W}(\mu_{-})$ and $\pi_{W}(\mu_{+})$. In particular, by choosing the subdivision constant to be $\sigma=\frac{1}{100}$ as in the statement of [Dur23, Proposition 4.7], we can produce $V_{i},V_{j}$ satisfying the desired conclusions.
@@
By [Raf14, Theorem 5.3], there is an active interval of times $I_{V_{i}}$ along the Teichmüller geodesic $\gamma$ during which the curves in $\partial V_{i}$ are shorter than $\epsilon$ and outside of which $\gamma$ has a bounded diameter projection to $\mathcal{C}(V_{i})$, and similarly for $V_{j},V_{k},W$. By our choice of the arrangement and location of $\partial V_{i},\partial V_{j},\partial V_{k}$ along any geodesic in $\mathcal{C}(W)$ between $\mu_{-},\mu_{+}$ and the fact that the projection of $\gamma$ to $\mathcal{C}(W)$ is a uniform (unparameterized) quasigeodesic [Raf14, Theorem 6.1], we may increase $L_{0}=L_{0}(S)>0$ as necessary to arrange that $I_{V_{i}},I_{V_{j}},I_{V_{k}}\subset I_{W}$ and that these intervals appear in this order along the parameter interval $[0,\infty)$ for $\gamma$.
@@
We also require some more organizational notation. Since the domains in $\mathcal{V}$ come in an order provided by the order of the slit curves $\zeta_{i}$, we can partition $\mathcal{V}$ into subcollections $\mathcal{V}_{j}$ of consecutive domains, with each of the #$\#V_{j}=N_{1}$ (for $N_{1}$ as in Claim 1), and the domains in $\mathcal{V}_{j}$ immediately preceding those in $\mathcal{V}_{j+1}$ for every $j\geq 1$.
@@
For any $W\in\mathcal{W}$ with $W\neq S$, we have #$\#J_{W}=1$.
@@
Claim 2 says that the $W_{i}$ are all distinct and, moreover, they have topological complexity at least $3$.
@@
Since items (1) and (2) follow from the generalities in [Dur23, Proposition 4.7], the only item to check is item (3). However, for each $i$, we have the containment $\zeta_{i},V_{i}\subset W_{i}$, so any container domain $S$ produced by [Dur23, Proposition 4.7] necessarily satisfies $\zeta_{j},V_{j},\subset W_{j}\subset X$ for any of the four indices produced in the claim. This completes the proof. ∎
Once again partitioning the domains in $\mathcal{W}$ as we did with $\mathcal{V}$ above into ordered subcollections, we can use Claim 3 to produce another group $\mathcal{Z}$ of container domains for the $W_{i}$ which are of complexity at least $4$. Moreover, it only takes $N_{1}$-many of the $W_{i}\in\mathcal{W}$ to produce each such container, while each such $W_{i}$ only took $N_{1}$-many of the $V_{j}\in\mathcal{V}$ to produce.
@@
We set some notation for the statement. Recall that Masur–Minsky [MM99] proved that Teichmüller geodesics project to unparameterized quasi-geodesics with uniform constants controlled by (the topology of) $S$. Since $\gamma$ has a minimal non-uniquely ergodic vertical foliation, its projection $\pi_{S}(\gamma)\subset\mathcal{C}(S)$ is a quasi-geodesic ray. Fix a geodesic ray $\Gamma$ in its asymptotic class [^2] Despite the fact that $\mathcal{C}(S)$ is locally infinite, one can find geodesic ray representatives using the finiteness properties of tight geodesics [Bow08].. For each $k$, let
@@
denote the closest-point projection to $\Gamma$ in $\mathcal{C}(S)$ of the projection to $\mathcal{C}(S)$ of the restriction of $\gamma$ to $[t_{k},t_{k+1}]$.
@@
    If $j=k-1$ or $j=k+1$ then $\mathrm{diam}_{\mathcal{C}(S)}(\Gamma_{k}\cap\Gamma_{j})<L/10$;
@@
For item (4), since $\pi_{S}(\gamma)$ is a uniform (unparameterized) quasi-geodesic, it can backtrack at most a bounded amount. By hyperbolicity of $\mathcal{C}(S)$, this backtracking is efficiently recorded on the geodesic ray $\Gamma$. In particular, by choosing $L=L(S)>0$ sufficiently large, we can guarantee that the overlap of $\Gamma_{k}$ with $\Gamma_{k-1}$ or $\Gamma_{k+1}$ is a controlled fraction of $L$.
@@ 5. Confirming sublinear Morseness
In this final section, we confirm that our examples are sublinearly Morse. For our purposes, the definition of sublinearly Morse is not strictly necessary; see [QRT24, DZ22]. Instead, we use the following criterion from [DZ22] for a Teichmüller geodesic to be sublinearly-Morse.
@@
As in Lemma 2 above, let $\Gamma$ be a geodesic representative in $\mathcal{C}(S)$ of the quasi-geodesic ray $\pi_{S}(\gamma)$. By the Bounded Geodesic Image Theorem [MM00, Theorem 3.1], if $\partial Y$ is not close to $\Gamma$, then $\gamma$ has bounded projections to $\mathcal{C}(Y)$.
On the other hand, the closest point projection $P_{Y}$ of $\partial Y$ to $\Gamma$ has uniformly bounded diameter. By increasing the constant $L$ in Lemma 2 a bounded amount as necessary, we can guarantee that $P_{Y}$ overlaps at most two consecutive $\Gamma_{k}$, and hence that $\partial Y$ is as far as we would like from the restrictions $\gamma_{k}=\pi_{S}(\gamma|_{[t_{k},t_{k+1}]})$ for all but two consecutive $k$.
In particular, this forces the active interval $I_{Y}$ to overlap at most two consecutive $[t_{k},t_{k+1}]$, while on the other hand the projection to $\mathcal{C}(Y)$ is coarsely constant outside of the active interval $I_{Y}$ for $Y$ by [Raf14, Theorem 6.1].
Supposing that $I_{Y}$ is contained in the union of $[t_{k-1},t_{k}]\cup[t_{k},t_{k+1}]$, then the distance formula for $\mathrm{Teich}(S)$ [Raf07, Dur16] and [Raf14, Theorem 6.1] provide that:
@@
(3) $\prec$ $\sum_{W}d_{W}(\gamma(t_{k-1}),\gamma(t_{k+1}))$
@@
In these computations, the coarse (in)equalities $\asymp$ and $\prec$ are (in)equalities which hold up to bounded additive and multiplicative constants depending only on the topology of $S$. Moreover, when $Y$ is an annulus, then the distance appearing in both lines (1) and (2) and the corresponding term in line (3) are all distances in the combinatorial horoball over $\mathcal{C}(Y)$ (which takes the length of the core curve of $Y$ into account), but this does not affect the computation. Finally, the penultimate inequality follows from item (1) of Lemma 2, where we replace $t_{k}$ with $t_{k-1}$ in line (5) depending (respectively) on whether $t\geq t_{k}$ or $t<t_{k}$.

### v2

@@ 4. Combinatorics of the slit curves
The proof involves an iterated argument which a blend of some results connecting short curves and subsurface projections due to Rafi [Raf05] and so-called “passing-up” techniques for producing large subsurface projections satisfying certain properties due to Durham [Dur23].
@@
We note one of the main technical difficulties here is that the slit curves $\zeta_{k}$ need not correspond to boundary curves of subsurfaces with large projections for $\mu^{-},\mu^{+}$. In particular, knowing that they are pairwise filling (Proposition 2) alone is not enough to show that they spread out in $\mathcal{C}(S)$ at a uniform rate.
Remark 4.
In the remaining proofs in this paper, when we write that some quantity is “uniform” or “uniformly bounded”, we mean that the quantity in question is controlled by the topology of $S$.
In the proof of Proposition 3, we will frequently use the following results due to Rafi [Raf05] and Durham [Dur23], which we include here for clarity. Please see [MM99b, MM00, Raf05, Raf07, Dur16] for background on subsurface projections and Teichmüller geometry.
The first statement we need is [Raf05, Theorem 6.1]. It says that short curves along Teichmüller geodesics determine large subsurface projections in their complement, and vice versa. The version below is quantitative—the shorter the curve, the larger the projection—and follows from the Rafi’s original proof in [Raf05], though it is not commonly stated this way in the literature. We need this quantitative version to know that the increasingly shorter slit curves produce increasingly larger projections, which we can then plug into the hierarchical machinery we will describe next.
Proposition 4 (Short curves and big projections).
Given a finite-type surface $S$, there exists constants $\epsilon_{0}=\epsilon_{0}(S)>0$ and $K_{0}=K_{0}(S)>0$ so that for any Teichmüller geodesic $(g,\mu^{-},\mu^{+})$ in $\mathrm{Teich}(S)$, the following hold:
  1. (1)
    For any $K\geq K_{0}$, there exists $\epsilon_{K}>0$ so that if $\alpha$ is a simple closed curve on $S$ so that $l_{g(t)}(\alpha)<\epsilon_{K}$ for some time $t$, then there exists a subsurface $Y\subset S-\alpha$ with $d_{Y}(\mu^{-},\mu^{+})>K$ (or $\log d_{Y}(\mu^{-},\mu^{+})>K$ in case $Y=Y_{\alpha}$ is the annulus about $\alpha$).
  2. (2)
    For any $\epsilon\geq\epsilon_{0}$, there exists $K_{\epsilon}>0$ so that if $d_{Y}(\mu^{-},\mu^{+})>K_{\epsilon}$ (or $\log d_{Y}(\mu^{-},\mu^{+})>K_{\epsilon}$ when $Y$ is an annulus), then there exists a time $t$ and a simple closed curve disjoint from $Y$ so that $l_{g(t)}(\alpha)<\epsilon$ for some time $t$.
The second statement we need is [Dur23, Proposition 4.7]. Roughly, it states that the boundary curves of a large collection $\mathcal{U}$ of subsurfaces to which a Teichmüller geodesic $g$ has a large projection $\pi_{W}(g)$ must spread out along the projection of $g$ to the curve graph $\mathcal{C}(W)$ of some subsurface $W\subseteq S$. Moreover, this statement is effective—the larger the cardinality of $\mathcal{U}$, the more of $\pi_{W}(g)$ they cover. We need it to gain quantitative control over how quickly the slit curves in our example spread out in the curve graph of the whole surface, $\mathcal{C}(S)$.
We need some terminology to set up the statement. Given a Teichmüller geodesic $(g,\mu^{-},\mu^{+})$ and a constant $K>0$, we say that $Y\subset S$ is $K$-relevant if $d_{Y}(\mu^{-},\mu^{+})>K$. We let $\mathrm{Rel}_{K}(\mu^{-},\mu^{+})$ denote the set of $K$-relevant domains for $g$.
We also need the notion of a $\sigma$-subdivision, which will allow us to precisely define what it means for boundary curves to spread out along the projection of a Teichmüller geodesic.
Definition 1 ($\sigma$-subdivision).
Let $W\subseteq S$ and $\gamma:I\to\mathcal{C}(W)$ be a geodesic path (i.e., segment, ray, or bi-infinite) in $\mathcal{C}(W)$ between $\pi_{W}(\mu^{-}),\pi_{W}(\mu^{+})$ coming from a geodesic $(g,\mu^{-},\mu^{+})$. For $\sigma>0$, we say that a subdivision $\{x_{i}\}$ of $I$ is a $\sigma$-subdivision of $\gamma$ if the $x_{i}$ decompose $I$ into a collection of subintervals $x_{i},x_{i+1}]$ so that for all but at most one $i$ we have $|x_{i+1}-x_{i}|=\sigma$, with the (possibly nonexistent) extra subinterval for which $|x_{i+1}-x_{i}|<\sigma$.
Given a $\sigma$-subdivision $\{x_{i}\}$ of a geodesic $\gamma$ between $\pi_{W}(\mu^{-}),\pi_{W}(\mu^{+})$ in $\mathcal{C}(W)$ for $W\subseteq S$ and a collection of relevant subsurfaces $\mathcal{V}$ of $W$, we let $\mathcal{V}_{i}$ denote the set of domains $V\in\mathcal{V}$ so that $p_{\gamma}(\partial V)\cap[x_{i},x_{i+1}]\neq\emptyset$, where $p_{\gamma}$ denotes closest point projection in $\mathcal{C}(W)$ to $\gamma$. The following proposition is [Dur23, Proposition 4.7] specialized to our setting:
Proposition 5 (Strong passing-up).
There exists $E=E(S)>0$ so that for any $K_{2}\geq K_{1}\geq 50E$, there exists $P_{1}=P_{1}(K_{1},K_{2})>0$ so that if $(g,\mu^{-},\mu^{+})$ is a Teichmüller geodesic in $\mathrm{Teich}(S)$, the following hold. If $\mathcal{V}\subset\mathrm{Rel}_{K_{1}}(\mu^{-},\mu^{+})$ with $\mathcal{V}>P_{1}$, then there exists $W\in\mathrm{Rel}_{K_{2}}(\mu^{-},\mu^{+})$ and $\mathcal{V}^{\prime}\subset\mathcal{V}$ so that $V\sqsubset W$ for all $V\in\mathcal{V}^{\prime}$ and
$\mathrm{diam}_{W}\left(\bigcup_{V\in\mathcal{V}^{\prime}}\rho^{V}_{W}\right)>K_{2}.$
Moreover, for any $\sigma\geq 10E$ and $n\in\mathbb{N}$, there exists $P_{2}(K_{1},K_{2},\sigma,n)>0$ so that if $\mathcal{V}>P_{2}$, then we can arrange the following to hold:
  • If $\gamma$ is any geodesic in $\mathcal{C}(W)$ between $\pi_{W}(\mu^{-}),\pi_{W}(\mu^{+})$ and $\{x_{i}\}$ is a $\sigma$-subdivision of $\gamma$ determining sets $\mathcal{V}_{i}$ as before, then
    $\{i|\mathcal{V}_{i}\neq\emptyset\}\geq n.$
Remark 5.
The original version of [Dur23, Proposition 4.7] is proven is context of hierarchically hyperbolic spaces. The Teichmüller space of any finite-type surfaces with the Teichmüller metric is an HHS, see [Raf07, Dur16, BHS19].
@@
By Proposition 2, we may assume that the slit curves $\zeta_{k}$ pairwise fill $S$. We fix a subsurface projection threshold constant $L_{0}=L_{0}(S)>0$ to be sufficiently large to make the arguments below work.
For sufficiently large $L_{0}=L_{0}(S)>0$, we can fix $\epsilon=\epsilon(L_{0},S)>0$ to be small enough so that if $\sigma$ is a simple closed curve on $S$ so that $l_{\gamma(t)}(\sigma)<\epsilon$ for some time $t$, then item (1) of Proposition 4 provides a subsurface $V\subset S-\zeta$ (with possibly $V$ the annulus with core $\zeta$) so that $\mathrm{diam}_{\mathcal{C}(V)}(\gamma)>L_{0}$.
@@
Consider the sequence of slit curves $\zeta_{k}$. By item (3) of Proposition 1, there exists $K_{0}=K_{0}(\epsilon)>0$ so that if $k>K_{0}$, then $l_{\gamma(t_{k})}(\zeta_{k})<\epsilon$. For each $k>K_{0}$, let $V_{k}$ denote the $L_{0}$-large subsurface in the complement $S-\zeta_{k}$ provided by item (2) of Proposition 4. We denote the collection of these subsurfaces $V_{k}$ by $\mathcal{V}$.
Our goal is to show that the $\zeta_{k}$ spread out in $\mathcal{C}(S)$ uniformly quickly. Using the $V_{k}$ as surrogates, the following claim forces the boundary curves $\partial V_{k}$ of any sufficiently large subcollection of the $V_{k}$ to spread out in some subsurface, with extra control over where the slit curves lie:
@@
For any $L_{1}>L_{0}$, there exists $N_{1}=N_{1}(S,L_{0},L_{1})>0$ so that for any collection $\mathcal{U}\subset\mathcal{V}$ with $\mathcal{U}\geq N_{1}$, there exists a subsurface $W$ and $V_{j},V_{j},V_{k},V_{l}\subset W$ with $V_{i},V_{j},V_{k},V_{l}\in\mathcal{U}$ so that
@@
    All pairwise distances in $\mathcal{C}(W)$ between $\mu^{-},\mu^{+},\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ are larger than $\frac{L_{1}}{100}$, and $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$ appear in that order along any geodesic in $\mathcal{C}(W)$ between $\mu^{-},\mu^{+}$ (see Figure 3).
@@
$\mu^{-}$$\mu^{+}$$\mathcal{C}(W)$$\partial V_{i}$$\partial V_{j}$$\partial V_{k}$$\partial V_{l}$$\zeta_{i}$$\zeta_{j}$$\zeta_{k}$$\zeta_{l}$$\frac{L_{1}}{100}>$$\frac{L_{1}}{100}>$$\frac{L_{1}}{100}>$$\frac{L_{1}}{100}>$$\frac{L_{1}}{100}$
Figure 3. Using Proposition 5, we can arrange for the boundary curves $\partial V_{i},\partial V_{j},\partial V_{k},\partial V_{l}$, and hence $\zeta_{i},\zeta_{j},\zeta_{k},\zeta_{l}$, appear in order along any geodesic in $\mathcal{C}(W)$ between $\mu^{-},\mu^{+}$.
@@
Items (1) and (2) follows immediately from Proposition 5. In particular, by choosing the subdivision constant to be $\sigma=\frac{1}{100}$ as in the statement, we can produce $V_{i},V_{j}$ satisfying the desired conclusions.
@@
By [Raf14, Theorem 5.3], there is an active interval of times $I_{V_{i}}$ along the Teichmüller geodesic $\gamma$ during which the curves in $\partial V_{i}$ are shorter than $\epsilon$ and outside of which $\gamma$ has a bounded diameter projection to $\mathcal{C}(V_{i})$, and similarly for $V_{j},V_{k},W$. By our choice of the arrangement and location of $\partial V_{i},\partial V_{j},\partial V_{k}$ along any geodesic in $\mathcal{C}(W)$ between $\mu^{-},\mu^{+}$ and the fact that the projection of $\gamma$ to $\mathcal{C}(W)$ is a uniform (unparameterized) quasigeodesic [Raf14, Theorem 6.1], we may increase $L_{0}=L_{0}(S)>0$ as necessary to arrange that $I_{V_{i}},I_{V_{j}},I_{V_{k}}\subset I_{W}$ and that these intervals appear in this order along the parameter interval $0,\infty)$ for $\gamma$.
@@
We also require some more organizational notation. Since the domains in $\mathcal{V}$ come in an order provided by the order of the slit curves $\zeta_{i}$, we can partition $\mathcal{V}$ into subcollections $\mathcal{V}_{j}$ of consecutive domains, with each of the $V_{j}=N_{1}$ (for $N_{1}$ as in Claim 1), and the domains in $\mathcal{V}_{j}$ immediately preceding those in $\mathcal{V}_{j+1}$ for every $j\geq 1$.
@@
For any $W\in\mathcal{W}$ with $W\neq S$, we have $J_{W}=1$.
@@
Claim 2 says that the $W_{i}$ are all distinct. Moreover, the $W_{i}$ all have topological complexity at least $3$ since each properly contains a subsurface of complexity at least $2$.
@@
Since items (1) and (2) follow from Proposition 5, the only item to check is item (3). However, for each $i$, we have the containments $\zeta_{i},V_{i}\subset W_{i}$, so any container domain $Y$ produced by Proposition 5 necessarily satisfies $\zeta_{j},V_{j},\subset W_{j}\subset Y$ for any of the four indices produced in the claim. This completes the proof. ∎
Once again partitioning the domains in $\mathcal{W}$ as we did with $\mathcal{V}$ above into ordered subcollections, we can use Claim 3 to produce another collection $\mathcal{Z}$ of container domains for the $W_{i}$ which are of complexity at least $4$. Moreover, it only takes $N_{1}$-many of the $W_{i}\in\mathcal{W}$ to produce each such container, while each such $W_{i}$ only took $N_{1}$-many of the $V_{j}\in\mathcal{V}$ to produce.
@@
We set some notation for the statement. Recall that Masur–Minsky [MM99a] proved that Teichmüller geodesics project to unparameterized quasi-geodesics with uniform constants controlled by (the topology of) $S$. Since $\gamma$ has a minimal non-uniquely ergodic vertical foliation, its projection $\pi_{S}(\gamma)\subset\mathcal{C}(S)$ is a quasi-geodesic ray. Fix a geodesic ray $\Gamma$ in its asymptotic class [^1] Despite the fact that $\mathcal{C}(S)$ is locally infinite, one can find geodesic ray representatives using the finiteness properties of tight geodesics [Bow08].. For each $k$, let
@@
denote the closest-point projection to $\Gamma$ in $\mathcal{C}(S)$ of the projection to $\mathcal{C}(S)$ of the restriction of $\gamma$ to $t_{k},t_{k+1}]$.
@@
    If j=k1j=k-1 or $j=k+1$ then $\mathrm{diam}_{\mathcal{C}(S)}(\Gamma_{k}\cap\Gamma_{j})<L/10$;
@@
For item (4), since $\pi_{S}(\gamma)$ is a uniform (unparameterized) quasi-geodesic, it can backtrack at most a bounded amount. By hyperbolicity of $\mathcal{C}(S)$, this backtracking is efficiently recorded on the geodesic ray $\Gamma$. In particular, by choosing $L=L(S)>0$ sufficiently large, we can guarantee that the overlap of $\Gamma_{k}$ with $\Gamma_{k-1}$ or $\Gamma_{k+1}$ is a controlled fraction of $L$, e.g. $\frac{L}{10}$.
@@ 5. Confirming sublinear Morseness
In this final section, we confirm that our examples are sublinearly Morse. For our purposes, the definition of sublinearly Morse is not strictly necessary, though we include it below; see [QRT24, DZ22] for more background in our setting.
Just for the sake of completeness, we recall the definition of sublinearly Morseness in [QRT24, Definition 3.2], also see [DZ22, Definition 2.3]:
A sublinearly Morse geodesic ray $\gamma$ satisfies a sublinear version of the Morse property satisfied by geodesics in hyperbolic spaces. Roughly speaking, this says that quasigeodesics that start and end on $\gamma$ must stay sublinearly close, for some uniform sublinear function.
Let $\kappa:[0,\infty)\to[1,\infty)$ be a sublinear function, i.e. $\lim_{t\rightarrow\infty}\kappa(t)/t=0$. Let $X$ be a proper geodesic metric space and fix a basepoint $\mathfrak{o}\in X$.
Definition 2.
A closed set $Z\subseteq X$ is called $\kappa$-Morse if there is a proper function $m_{Z}:\mathbb{R}^{+}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$, such that for any sublinear function $\kappa^{\prime}$, any $r>0$, there exists $R\geq r$, such that for any $(q,Q)$-quasi-geodesic ray $\beta$ with $\beta(0)=\mathfrak{o}$, and $m_{Z}(q,Q)$ sufficiently small compared to $r$, we have
$d_{X}(\beta(R),Z)\leq\kappa^{\prime}(R)\implies\text{ for any }t\in[0,r],d(\beta(t),Z)\leq m_{Z}(q,Q)\kappa(d_{X}(\mathfrak{o},\beta(t)))$
  • The function $m_{Z}$ is called a Morse gauge of $Z$.
The following is a useful criterion from [DZ22] for a Teichmüller geodesic to be sublinearly-Morse.
@@
As in Lemma 3 above, let $\Gamma$ be a geodesic representative in $\mathcal{C}(S)$ of the quasi-geodesic ray $\pi_{S}(\gamma)$. By the Bounded Geodesic Image Theorem [MM00, Theorem 3.1], if $\partial Y$ is not uniformly close to $\Gamma$, then $\gamma$ has a uniformly bounded projection to $\mathcal{C}(Y)$.
On the other hand, the closest point projection $P_{Y}$ of $\partial Y$ to $\Gamma$ has uniformly bounded diameter. By increasing the constant $L$ in Lemma 3 a bounded amount as necessary (which Proposition 3 allows us to do), we can guarantee that $P_{Y}$ overlaps at most two consecutive $\Gamma_{k}$, and hence that $\partial Y$ is as far as we would like from the restrictions $\gamma_{k}=\pi_{S}(\gamma|_{[t_{k},t_{k+1}]})$ for all but two consecutive $k$.
In particular, this forces the active interval $I_{Y}$ to overlap at most two consecutive $t_{k},t_{k+1}]$, while on the other hand the projection to $\mathcal{C}(Y)$ is coarsely constant outside of the active interval $I_{Y}$ for $Y$ by [Raf14, Theorem 6.1].
Supposing that $I_{Y}$ is contained in the union of $t_{k-1},t_{k}]\cup[t_{k},t_{k+1}]$, then the distance formula for $\mathrm{Teich}(S)$ [Raf07, Dur16] and [Raf14, Theorem 6.1] provide that:
@@
(3) $\prec$ $\sum_{W\subset\mathcal{L}_{k}}d_{W}(\gamma(t_{k-1}),\gamma(t_{k+1}))$
@@
In these computations:
  • The coarse (in)equalities $\asymp$ and $\prec$ are (in)equalities which hold up to bounded additive and multiplicative constants depending only on the topology of $S$.
  • Given an annulus $V\subset S$, the distance $d_{V}(-,-)$ denotes distance in a combinatorial horoball over the annular curve graph for $V$ [Dur16] or alternatively an appropriate horoball in $\mathbb{H}^{2}$ [DDM14].
  • In line (3),
    $\mathcal{L}_{k}=\{W\subset S|d_{W}(\gamma(t_{k-1}),\gamma(t_{k+1})>M\}$
    denotes the subsurfaces where $\gamma(t_{k-1}),\gamma(t_{k+1})$ have a projection larger than $M=M(S)$, a constant depending only on the topology of $S$.
  • The transition from line (3) to (4) follows from the distance formula for the Teichmüller metric [Raf07] (see [Dur16, Theorem 7.14] or [DDM14, Proposition A.1] for the version used here).
  • Finally, the penultimate inequality follows from item (1) of Lemma 3, where we replace $t_{k}$ with $t_{k-1}$ in line (6) depending (respectively) on whether $t\geq t_{k}$ or $t<t_{k}$.